# Language Modeling for the Future of Finance: A Quantitative Survey into Metrics, Tasks, and Data Opportunities

### v1

@@ Abstract
Recent advances in language modeling have led to growing interest in applying Natural Language Processing (NLP) techniques to financial problems, enabling new approaches to analysis and decision-making. To systematically examine this trend, we review 374 NLP research papers published between 2017 and 2024 across 38 conferences and workshops, with a focused analysis of 221 papers that directly address finance-related tasks. We evaluate these papers across 11 qualitative and quantitative dimensions, identifying key trends such as the increasing use of general-purpose language models, steady progress in sentiment analysis and information extraction, and emerging efforts around explainability and privacy-preserving methods. We also discuss the use of evaluation metrics, highlighting the importance of domain-specific ones to complement standard machine learning metrics. Our findings emphasize the need for more accessible, adaptive datasets and highlight the significance of incorporating financial crisis periods to strengthen model robustness under real-world conditions. This survey provides a structured overview of NLP research applied to finance and offers practical insights for researchers and practitioners working at this intersection.
@@ 1 Introduction
Language modeling has become a key tool in natural language processing (NLP) for analyzing unstructured text, including reports, news articles, and social media posts. These methods are increasingly applied to finance-related tasks such as sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), summarization (Khanna et al., 2022), stock prediction (Jain & Agrawal, 2024), and volatility forecasting (Niu et al., 2023). Given the increasing number of NLP papers addressing financial problems, there is a need to systematically examine how the field is contributing to finance-related applications. While prior surveys have examined this intersection (Table 1), many adopt a qualitative approach (Chen et al., 2022b; Gao et al., 2021b; Xiao et al., 2024a), focusing on broad NLP techniques (Jagdale & Deshmukh, 2025; Man et al., 2019; Liu, 2024), deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023c), or specific tasks such as sentiment analysis (Mishev et al., 2020). However, these studies often lack systematic evaluation.
In this study, we focus specifically on NLP research applied to finance. Our scope includes papers published in NLP venues where methods are developed, benchmarked, or adapted for financial problems. We analyze 374 papers published from 2017 to 2024 across 38 NLP conferences and workshops. After further filtering (Section 2), we retain 221 papers that directly address finance-related tasks. These are evaluated across 11 qualitative and quantitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility.
Our analysis reveals several patterns in the current research landscape. Commonly addressed tasks include sentiment analysis, information extraction, and question answering, while areas such as explainability and privacy-preserving techniques are explored less frequently (Section 3). Standard machine learning metrics are widely used for evaluation, although these may not always reflect domain-specific priorities (Section 7.1). Most studies focus on stable market conditions, with less attention to more volatile periods (Section 7.2) and relying on outdated sources or overlooking issues such as survivorship bias (Section 7.3) Finally, we observe a trend toward the use of general-purpose language models over custom architecture (Section 4). While this reflects the rapid integration of recent advances in NLP, further exploration of domain-specific strategies may offer additional value for finance-related tasks.
@@
(Chen et al., 2020b) 2016-2019 62 3 Financial Technology × × × ×
(Chen et al., 2022a) 2018-2022 38 2 Financial Technology × × ×
@@
(Zhao et al., 2024a) 2004 (A)-2024 146 1 Large Language Models × × ×
@@
Table 1: Comparison of previous Natural Language Processing surveys in finance based on their focus areas, number of papers reviewed, and analytical features. Due to the absence of systematic collection methods in most prior research work, entries marked with (A) in the "Review Years" column indicate that the reported range is approximate.
@@ 2 Paper Extraction Process
Figure 1: Overview of our paper selection process and analysis dimensions. We collected papers from a broad range of NLP venues using abstract-level keyword filtering, yielding 374 candidates. After removing mismatches and shared task papers, we retained 221 papers, categorized into four groups based on their connection to financial tasks. The analysis focuses on the selected dimensions as shown in (Table 5, Appendix B).
To study how NLP has been applied to finance, we selected papers from 38 NLP venues, including CoLM, ACL, NAACL, EMNLP, LREC, COLING, and workshops like FinNLP and the Workshop on Economics and Natural Language Processing. The full list of venues is provided in Table 9 (Appendix E). As illustrated in Figure 1, we began by filtering papers that mentioned “finance” or “financial” in their abstracts. Out of all such papers published from 1975 to the present, more than 94% were published from 2017 onward. This growth coincides with the emergence of transformer-based models (Vaswani et al., 2017), which greatly expanded the scope of NLP applications. Accordingly, we use 2017 as a threshold year for our analysis.
@@
This initial filtering returned 374 papers. Among them, 88 were tied to shared tasks (e.g., SemEval-2017 (Kar et al., 2017), FinCausal-2022 (Mondal et al., 2022)), and 65 used the term “financial” in a different contexts – such as referencing financial resources (Sekeres et al., 2024; Ding & Riloff, 2018) – without any actual financial application. After removing these, we retained 221 papers for our analysis.
Unlike previous surveys, which often emphasize qualitative observations, our study combines both qualitative and quantitative methods to provide a broader view of the field. We categorized the selected papers by their primary focus, resulting in four task-based categories (Figure 2). This classification helps us examine trends in how NLP techniques are being used to tackle finance-relevant problems and supports more detailed comparisons across types of tasks.
3 Task Distribution in NLP for Financial Applications
@@
To understand how NLP has been used in financial contexts, we categorized papers into four groups (shown in Table 2) based on their primary tasks (Figure 2). This classification helps map how language modeling techniques are applied and how directly each paper engages with finance, with Category I representing the most direct financial use cases. Detailed descriptions of each category are provided in Appendix D.
Financial Forecast (Category I) papers cover predictive tasks such as stock price and volatility forecasting. While these areas are well-studied, others like economic forecasting (Arno et al., 2023), risk assessment (Zhou et al., 2020), and cryptocurrency prediction (Seroyizhko et al., 2022) receive less focus. These areas present opportunities for expanding the reach of prediction models and improving their robustness and comparability (Section 7).
Financial Resources (Category II) papers often focus on dataset construction (Section 4, Figure 3). These datasets – ranging from annotated earnings calls to news and speeches – support tasks like financial event extraction (Huang et al., 2024; Ju et al., 2023), fraud detection (Erben & Waldis, 2024; Wang et al., 2019), and annotation (Aguda et al., 2024; Khatuya et al., 2024). Although these tasks contribute to enhancing financial data processing, other tasks remain relatively underrepresented.
In the Financial Applications (Category III) group, sentiment analysis (Rodriguez Inserte et al., 2023), information extraction (Lior et al., 2024), and question answering (Kosireddy et al., 2024) are the dominant tasks. These methods are widely used to analyze earnings calls, reports, and market commentary, where investor sentiment and factual extraction are key inputs for decision-making. Recent efforts have aimed to improve QA models for handling financial text (Theuma & Shareghi, 2024; Liu et al., 2024; Mavi et al., 2023).
Finally, among Finance-Related (Category IV) papers, the most studied areas are dataset construction and numerical reasoning. While these datasets are not finance-specific, they have potential applications in finance. For instance, fake news detection datasets (Vargas et al., 2021) can help reduce misinformation in markets. Numerical reasoning (Akhtar et al., 2023), important for understanding financial statements and assessing risk, is receiving more attention. However, other areas – such as explainable AI (XAI) (Klein & Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) – remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
4 Language Model Adoption & Datasets
Figure 3: Timeline of PLM/LLM adoption in NLP research applied to finance, alongside key datasets by task type: C (custom models), QA (question answering), IE (information extraction), OL (other language resources), and FD (financial documents). For each model, the top date indicates release; the bottom, first usage in the surveyed papers.
Acceleration of Model Adoption
Figure 3 (Language Models) shows how quickly language models are adopted for finance-related NLP tasks. While early models like BERT (Devlin et al., 2019) took 13 months to appear in these applications, recent models – GPT-3.5, GPT-4 (Achiam et al., 2023), and Mistral (Jiang et al., 2023) – were adopted within 6–12 months. LLaMA-2 (Touvron et al., 2023)) saw integration in just 5 months, reflecting the community’s increasing speed in applying general NLP advances to financial tasks.
Dominance of General-Purpose Models
As shown in Figure 3 (Language Models), general-purpose models – LLMs like GPT-4 and LLaMA-2, and Pretrained Language Models (PLMs) like BERT and RoBERTa (Liu et al., 2019)) – are used far more frequently than domain-adapted ones. BERT appears in over 70 papers, while finance-tuned models such as FinBERT (Araci, 2019; Yang et al., 2020b; Liu et al., 2020; Huang et al., 2022), FinMA (Xie et al., 2024a), and FinGPT (Yang et al., 2023) are less common. Notably, earlier FinBERT versions are more widely used than later ones, suggesting that improvements in general models often outweigh the benefits of domain-specific adaptation.
Decline of Custom Architectures
Custom models were common before 2022, especially for tasks like table-based QA (e.g., TagOp for TAT-QA (Zhu et al., 2021), HyBrider for HybridQA (Chen et al., 2020c)). Some models, like MDRM (Qin & Yang, 2019), HTML (Yang et al., 2020a), and HAN (Hu et al., 2021), are still used as baselines, particularly in volatility prediction (Niu et al., 2023; Mathur et al., 2022). However, the rise of general LLMs has led to a sharp decline in custom architectures. As a result of the accelerated development and deployment of these general purpose models, researchers increasingly rely on adapting pretrained models instead of building domain-specific ones from scratch.
Language Coverage in Datasets
The OL and FD sections of Figure 3 highlights the scarcity of non-English datasets for forecasting tasks such as return or volatility prediction. One exception is CMIN (Luo et al., 2023), which includes English and Chinese financial data. Most datasets use English-only sources like EDGAR filings (Maheshwari et al., 2022) and financial news (Xie et al., 2024b), which may limit model performance in non-English markets. Creating multilingual forecasting datasets would support broader applicability and reduce language bias, as also noted by Jørgensen et al. (2023).
Since 2017, research applying NLP to finance has shifted from custom models toward general-purpose pretrained architectures. New models are adopted faster than before, and general PLMs have largely replaced finance-specific ones. This preference also influences dataset use: most forecasting datasets are English-only, limiting generalization to global markets.
5 Methods Used in NLP for Finance
Figure 4: Trends in methodology adoption over time, showing stable use of traditional techniques and rapid growth in PLMs and LLMs. Less frequent methods are omitted for clarity.
To understand the methodological landscape, we grouped the approaches used in the surveyed papers into nine categories (Appendix C, Table 6). While the total number of papers has grown, methods like statistical NLP, conventional machine learning, and embeddings have remained mostly flat. In contrast, there has been a sharp increase in the use of deep learning, PLMs and LLMs, reflecting the field’s shift toward pretrained architectures.
Shift Toward PLMs and LLMs
As seen in Figure 4, the first three categories – statistical NLP, embeddings, and conventional ML – have persisted but have seen reduced usage as PLMs and LLMs take lead. The rise began with models like BERT in 2019, which significantly improved context handling in financial texts (Li et al., 2017; Chen et al., 2019; Ein-Dor et al., 2019). PLMs are now widely used for tasks such as summarization and entity recognition. Since 2023, LLMs like ChatGPT-3.5 have seen rapid adoption (Tang et al., 2023; Li et al., 2023a; Ni et al., 2023), especially for text generation (Guo et al., 2023b) and financial QA (Uluoglakci & Temizel, 2024). This transition has also brought more complex deep learning architectures used for multi-step reasoning and domain-specific challenges (Ma et al., 2024; Zhao et al., 2024b; Su et al., 2024).
Underexplored Techniques
While pretrained models dominate, several methods remain underused despite their promise. Statistical modeling, graph-based approaches, and knowledge graphs could improve model robustness and interpretability, particularly for applications like causality detection (Tabari et al., 2018) or market dynamics (Menzio et al., 2024). Emerging techniques like GraphRAG (Edge et al., 2024) highlight the potential of such methods, yet they appear in only a fraction of papers. Their limited presence suggests opportunities for expanding methodological diversity in future research.
6 Code & Data Accessibility
Figure 5: Temporal trends in code and dataset availability, highlighting the shift toward open-source practices and the growing accessibility of NLP resources for finance.
As the adoption of language models accelerates, the issue of open accessibility in NLP research, especially in finance, becomes more relevant. As shown in Figure 5, open-source practices have become more common in recent years. Before 2021, closed-source code dominated, but the trend has shifted toward sharing code and datasets. This move enhances transparency, reproducibility, and collaborative progress in NLP research applied to finance (Whited, 2023). Despite this, some papers still include inactive links or empty repositories, limiting reproducibility. At the same time, dataset availability has steadily increased since 2017, pointing to growing awareness around open data. Maintaining functional, up-to-date repositories remains essential to support meaningful benchmarking and model development.

### v2

# Language Modeling for the Future of Finance: A Survey into Metrics, Tasks, and Data Opportunities

@@ Abstract
Recent advances in language modeling have led to growing interest in applying Natural Language Processing (NLP) techniques to financial problems, enabling new approaches to analysis and decision-making. To systematically examine this trend, we review 374 NLP research papers published between 2017 and 2024 across 38 conferences and workshops, with a focused analysis of 221 papers that directly address finance-related tasks. We evaluate these papers across 11 quantitative and qualitative dimensions, and our study identifies the following opportunities: (i) expanding the scope of forecasting tasks; (ii) enriching evaluation with financial metrics; (iii) leveraging multilingual and crisis-period datasets; and (iv) balancing PLMs with efficient or interpretable alternatives. We identify actionable directions for research and practice, supported by dataset and tool recommendations, with implications for both the academia and industry communities.
@@ 1 Introduction
Language modeling has become a key tool in natural language processing (NLP) for analyzing unstructured text, including reports, news articles, and social media posts. These methods are increasingly applied to finance-related tasks such as sentiment analysis (Balakrishnan et al., 2022), information extraction (Huang et al., 2023), summarization (Khanna et al., 2022), stock prediction (Jain and Agrawal, 2024), and volatility forecasting (Niu et al., 2023). Given the increasing number of NLP papers addressing financial problems, there is a need to systematically examine how the field is contributing to finance-related applications. While prior surveys have examined this intersection (Table 1), many lack a structured approach (Chen et al., 2022a; Gao et al., 2021b; Xiao et al., 2024), focusing on broad NLP techniques (Jagdale and Deshmukh, 2025; Man et al., 2019; Liu, 2024), deep learning (Ozbayoglu et al., 2020) and Large Language Models (LLMs) (Nie et al., 2024; Li et al., 2023), or specific tasks such as sentiment analysis (Mishev et al., 2020).
In this study, we focus specifically on NLP research applied to finance. Our scope includes papers published in NLP venues where methods are developed, benchmarked, or adapted for financial problems. We analyze 374 papers published from 2017 to 2024 across 38 NLP conferences and workshops. After further filtering (Section 2), we retain 221 papers that directly address finance-related tasks. These are evaluated across 11 quantitative and qualitative dimensions, including tasks, methodologies, datasets, evaluation metrics, and accessibility.
Table 1: Comparison of previous Natural Language Processing surveys in finance based on their focus areas, number of papers reviewed, and analytical features. Due to the absence of systematic collection methods in most prior research work, entries marked with (A) in the "Review Years" column indicate the approximate range.
@@
(Chen et al., 2020a) 2016-2019 62 3 Financial Technology × × × ×
(Chen et al., 2022b) 2018-2022 38 2 Financial Technology × × ×
@@
(Zhao et al., 2024) 2004 (A)-2024 146 1 Large Language Models × × ×
@@
Our analysis reveals not only valuable insights but also actionable directions for both research and practice with dataset and tool recommendations. Forecasting tasks, while well-established, leave room for exploration, especially in areas such as risk and macroeconomic prediction (Section 3). Financial evaluation metrics are gaining traction and could further improve the practical applicability of models (Section 4.1). The growing availability of temporally diverse (Sections 4.2, 4.3), multilingual and multimodal (Sections 4.4) datasets enables more robust, globally applicable models. Finally, while PLMs and LLMs dominate, lightweight and interpretable methods remain viable for practical deployment (Section 5).
@@ 2 Paper Extraction Process
To study how NLP has been applied to finance, we selected papers from 38 NLP venues, including ACL, NAACL, EMNLP, LREC, CoLM, COLING, and workshops like FinNLP and the Workshop on Economics and Natural Language Processing. As illustrated in Figure 1, we began by filtering papers that mentioned “finance” or “financial” in their abstracts. Out of all such papers published from 1975 to the present, more than 94% were published from 2017 onward. This growth coincides with the emergence of transformer-based models (Vaswani et al., 2017), which greatly expanded the scope of NLP applications. Accordingly, we use 2017 as a threshold year for our analysis. The keyword filtering approach enabled us to cast a wide net, allowing for a high volume of papers. We were then able to filter these further using manual methods.
This initial filtering returned 374 papers. Among them, 88 were tied to shared tasks (e.g., SemEval-2017 (Kar et al., 2017), FinCausal-2022 (Mondal et al., 2022)), and 65 used the term “financial” in a different contexts (such as referencing financial resources (Sekeres et al., 2024; Ding and Riloff, 2018)) without any actual financial application. After removing these, we retained 221 papers for our analysis.
Unlike previous surveys, which often emphasize qualitative observations, our study combines quantitative and qualitative methods for a broader view of the field. We categorized the selected papers by their primary focus into four task-based categories (Figure 2). This classification helps us examine trends in how NLP techniques are used to tackle finance-relevant problems and supports more detailed comparisons across types of tasks.
3 Task Distribution in NLP for Financial Applications
To understand how NLP has been used in financial contexts, we categorized papers into four distinct groups (shown in Table 2) based on their primary tasks (Figure 2). This classification provides a mapping from language modeling techniques to their applications and the level of engagement with finance, with Category I papers representing direct financial use cases.
@@ 3 Task Distribution in NLP for Financial Applications
Figure 1: Overview of our paper selection process and analysis dimensions. We collected papers from a broad range of NLP venues using abstract-level keyword filtering, yielding 374 candidates. After removing mismatches and shared task papers, we retained 221 papers, categorized into four groups by their connection to financial tasks.
Financial Forecast (Category I) papers cover predictive tasks such as stock price and volatility forecasting. While these tasks are well-studied, areas like economic forecasting (Arno et al., 2023), risk assessment (Zhou et al., 2020), and cryptocurrency prediction (Seroyizhko et al., 2022) receive less focus. These areas offer opportunities to expand the reach of predictive models, improving robustness and comparability (Section 4).
Financial Resources (Category II) papers often focus on dataset construction (Figure 6). Ranging from annotated earnings calls to news and speeches, These datasets support tasks like financial event extraction (Huang et al., 2024; Ju et al., 2023), fraud detection (Erben and Waldis, 2024; Wang et al., 2019), and annotation (Aguda et al., 2024; Khatuya et al., 2024). While these tasks enhance financial data processing, other tasks remain relatively underrepresented.
In the Financial Applications (Category III) group, sentiment analysis (Rodriguez Inserte et al., 2023), information extraction (Lior et al., 2024), and question answering (Kosireddy et al., 2024) are the dominant tasks. These methods are widely used to analyze earnings calls, reports, and market commentary, where investor sentiment and factual extraction are key inputs for decision-making. Recent efforts have aimed to improve QA models for handling financial text (Theuma and Shareghi, 2024; Liu et al., 2024; Mavi et al., 2023).
Finally, among Finance-Related (Category IV) papers, the most studied areas are dataset construction and numerical reasoning. While these datasets are not finance-specific, they have potential applications in finance. For instance, fake news detection datasets (Vargas et al., 2021) can help reduce misinformation in markets. Numerical reasoning (Akhtar et al., 2023), important for understanding financial statements and assessing risk, is receiving more attention. However, other areas – such as explainable AI (XAI) (Klein and Walther, 2024) and privacy-preserving methods (Abbe et al., 2012) – remain rarely explored, despite their relevance to secure and interpretable decision-making (Basu et al., 2021).
@@ 4 Potential in Financial Forecasting
Category I papers reveal several areas where forecasting models could be improved. Most studies use general ML metrics, leaving out finance-specific measures that better capture practical utility (Section  4.1). Crisis periods are rarely considered, limiting insights into how models behave under stress (Section  4.2). Finally, there is a strong U.S. and English-language bias in data, with limited adoption of global financial datasets (Section  4.3,  4.4). Expanding these areas of exploration could strengthen model robustness and support broader real-world applicability.
4.1 Advancing Evaluation with Finance Specific Metrics
As shown in Figure 3, Category I papers mostly rely on ML metrics such as accuracy, F1, or MSE (Sawhney et al., 2020; Wu, 2020; Yangjia et al., 2022). While useful, these do not fully reflect financial performance. Finance-specific metrics offer risk-adjusted insights that ML measures overlook. Incorporating such metrics would improve comparability across models and align research outcomes with practical needs (Zhang et al., 2024; Zou et al., 2022; Sawhney et al., 2021).
Consider two models that predict daily stock returns. Model A achieves a lower mean squared error (MSE) than Model B, yet when deployed in a trading strategy, it results in frequent trades, high volatility, and poor performance during market downturns. In contrast, Model B, despite its higher MSE, generates a smoother equity curve, experiences lower drawdown, and delivers a higher Sharpe Ratio thus making it more appealing from an application perspective. This highlights a critical insight emphasized in financial machine learning literature: traditional statistical metrics like MSE are often inadequate for evaluating trading strategies, as they fail to capture real-world profitability and risk exposure (Bailey et al., 2015; de Prado, 2018). Instead, financial metrics such as Sharpe Ratio, Maximum Drawdown, and Cumulative Return are essential because they reflect the model’s risk-adjusted performance and robustness across market regimes (Lo, 2002; Krauss et al., 2017; Takahashi et al., 2009) as well as the applicability of the model to practical real world situations.
Key Financial Metrics
There are certain financial metrics that measure models’ performance in a financial setting and provide insight into the their applicability. Sharpe Ratio quantifies risk-adjusted return by comparing excess returns to volatility (Investopedia, 2025a). Maximum Drawdown (MDD) captures the largest loss from peak to trough, indicating downside risk and model robustness (Hayes, 2024). Cumulative Return measures total profit over a period: a direct benchmark for model outputs (Investopedia, 2007).
Tools for Metric Calculation
Several Python libraries support quantitative finance and model-driven investment strategies. PyPortfolioOpt (Martin, 2021) offers portfolio optimization tools including Sharpe Ratio and risk-return metrics. QuantLib (QuantLib Team, 2000) provides analytics for pricing, drawdown, and custom risk modeling. Backtrader (Backtrader Team, 2015) enables backtesting of trading strategies with performance evaluation, ideal for simulating NLP-driven trading. bt (Morissette, 2025) supports rapid prototyping and comparison of forecasting models with built-in financial metrics.
4.2 Strengthening Robustness by Incorporating Crisis Periods
Figure 4 shows that researchers use financial data dating back to 1993, when electronic filings became publicly available, with another rise in 2005 after the HTML filing mandate (U.S. SEC, 2025). However, most publications use post-2009 (the global financial crisis) and pre-2020 (the COVID-19 pandemic) data, thus missing the crisis periods in evaluation. Crisis periods are essential for evaluating model robustness, allowing stress testing (Investopedia, 2025b) and helping build forecasting models that remain reliable under instability.
Consider the case of Long-Term Capital Management (LTCM). Its quantitative models were highly leveraged and calibrated almost exclusively on stable market data, failing to account for extreme tail events. When the 1997 Asian and 1998 Russian crises struck, LTCM’s Value-at-Risk (VaR) models severely underestimated exposure, leading to large-scale losses as spreads exploded and liquidity evaporated. Within weeks, LTCM lost over 90% of its equity, triggering systemic risk concerns across the global financial system (Department of Land Economy, 2003; Federal Reserve History, 2025; Vanity Fair Staff, 1998). As noted in academic and regulatory analyses, this collapse illustrates the dangers of relying solely on backtests from stable periods without incorporating the possibility of rare but catastrophic shocks.
Most current models are trained in stable conditions that may not reflect their behavior during volatility. While there were data availability challenges prior to 2015 (Bloomberg L.P., 2015), the crisis years remain underused. In Section 4.3, we highlight valuable data sources, including those with crisis years data.
4.3 Opportunities in Data Coverage
Most forecasting models use common financial sources: stock prices, SEC filings, financial news, earnings calls. But many valuable datasets are either rarely used (e.g., Federal Reserve reports (Shah et al., 2023; Menzio et al., 2024)) or completely absent from NLP research (e.g., shareholder letters). Public resources like FRED (Federal Reserve Bank of St. Louis, 2024) and Fama-French (Fama and French, 2024) offer rich but underexplored macro indicators. Incorporating these datasets could help capture macroeconomic trends and market dynamics, improving model reliability. Below, we highlight sources of valuable financial data, including those already used by the community and the other beneficial ones.